The curve satisfying the eqn 


dy / dx = (y2 - 2xy - x2) / (y2 + 2xy - x2)    passes through (-1, 1) is a 


(a) staight line      (b) circle    (c) ellipse     (d) none of these

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit Gupta

dy / dx = (y2 - 2xy - x2) / (y2 + 2xy - x2)

Let y=vx

dy/dx =v+xdv/dx

 

so v+x dy/dx =(v2 - 2v -1) / (v2 + 2v - 1)

x dv/dx =(v2 - 2v -1) / (v2 + 2v - 1)     -v

x dv/dx = -(v3 +v2 +v +1)/(v2 + 2v - 1)

(v2 + 2v - 1)/(v3 +v2 +v +1)  dv =-dx/x

 

Now variable are seperated you can solve the differential equation


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin